# Collapsing in polygonal dynamics

### v2

@@ 2 Basic properties and examples
It is exactly same proof as in [7, Prop 3.2], which treated the case of $\mathbb{P}^{2}(\mathbb{C)}$. The deformation by the scaling symmetry takes place on $\mathcal{T}_{n}$, but we can still take a smooth lift on $\tilde{\mathcal{T}}_{n}$ which will commute with the dynamic. One example [^6] Which is the one we use in section 4. of such a lift is the family of representants $(P_{z})$ for which the first $d+1$ vertices $p_{1}(z),p_{2}(z),\dots,p_{d+1}(z)$ are fixed on
$1:\dots:1],[1:0:\dots:0],\dots,[0:\dots:0:1].$
Because of [14, Prop. 4.5.10], this is the unique family of representant of $([P]_{z})$ satisfying this property.
Now, let $P_{z},\penalty 10000\ \tilde{P}_{z}$ be two lifts such that $P_{1}=\tilde{P}_{1}$, and $\phi_{z}\in\mathbb{P}\mathrm{GL}_{d+1}$ be a smooth family such that $\tilde{P}_{z}=\phi_{z}\cdot P_{z}$. This forces $\phi_{1}=\mathrm{Id}$ and $\tilde{M}_{z}=\phi_{z}M_{z}\phi_{z}^{-1}$. By differentiating, which we can do since $k$ is a complete metric space, we get:
$\left.\frac{d\tilde{M}_{z}}{dz}\right|_{z=1}=\left.\frac{dM_{z}}{dz}\right|_{z=1}+\left[\left.\frac{d\phi_{z}}{dz}\right|_{z=1},M_{1}\right].$
Since $P_{1}$ is a closed polygon, we have $M_{1}=\mathrm{Id}$ and so the commutator vanishes. Hence infinitesimal monodromy is well defined. [^7] As it was remarked in [7], the infinitesimal monodromy is not well defined for twisted polygons, since the only central element of $\mathbb{P}\mathrm{GL}_{d+1}$ is $\mathrm{Id}$.
The infinitesimal monodromy is defined from the family of monodromies $(M_{z})_{z\in Z}$ coming from the scaling symmetry. But $T$ preserves the monodromy and commutes with the scaling symmetry, so it also preserves the infinitesimal monodromy. ∎
The rest of this section are tools for the collapse conjecture 3.2. Next lemma says that if $k$ is an algebraicly closed complete valued field, for a non-trivial valuation, then iterating a generic element of $\mathbb{P}\mathrm{GL}_{d+1}(k)$ yields a North-South dynamic.
@@
Let $k$ be an algebraicly closed, complete field for the metric coming from a non-trivial valuation.
Then a generic element of $\mathbb{P}\mathrm{GL}_{d+1}(k)$ has one attractive and one repelling fixed point.
Proof.
Let $M\in\mathrm{GL}_{d+1}(k)$. The set of diagonalisable matrices of $\mathrm{GL}_{d+1}(k)$ is Zariski dense; indeed, $M$ is diagonalisable if its characteristic polynomial $\chi_{M}(X)$ have distinct roots. This happens if its discriminant, which is a polynomial in the coefficients $\Delta(a_{1,1},\dots,a_{d+1,d+1})$, doesn’t vanish. Since this is not the zero polynomial, we obtain a Zariski-open set. It is not empty since it contains the identity.
@@
with $|\lambda_{1}|\leq\dots\leq|\lambda_{d+1}|$. Since the valuation is non-trivial, then out of genericity we can assume that the inequalities are in fact strict. In $\mathbb{P}\mathrm{GL}_{d+1}$, we can take the representant with $\lambda_{d+1}=1$. So:
@@
Thus $M$ has an attractive fixed point in $\mathbb{P}^{d}$, being $0:\dots:0:1]$ (the origin).
We can do similarly for $M^{-1}$ and take a representant with $\lambda_{1}=1$. This yields:
@@
The following fact is classical but we’ll prove it anyway. If we apply it in $\mathbb{P}\mathrm{GL}_{d+1}$, shared eigenlines translates to shared fixed points.
Lemma 2.3.
Let $M,M^{\prime}\in\mathrm{GL}_{d+1}$ be two commuting matrices such that all the eigenvalues of $M$ are distinct. Then $M$ and $M^{\prime}$ share the same eigenlines.
Proof.
Let $x$ be such that $Mx=\lambda x$. Hence we have:
$MM^{\prime}x=M^{\prime}Mx=\lambda M^{\prime}x,$
so $M^{\prime}x$ is an eigenvector of $M$ for the eigenvalue $\lambda$. Since the eigenvalues are all distinct, $M^{\prime}x$ and $x$ lie on the same eigenline of $M$, and thus $x$ is an eigenvector of $M^{\prime}$. ∎
Let’s finish this section with a definition of discrete integrability, inspired by [1]. It aims to retrieve the essential property of the various flavours of integrability: the Liouville-Arnold one (involving symplectic forms, hamiltonians and lagrangian toris) or the algebraic one (involving Lax pairs, theta functions and abelian varieties) [^8] See [15] for a review of integrable systems.. The polygonal dynamics we consider are usually integrable on their moduli space.
Definition 2.4.
A dynamical system $(X,T)$ is said to be discretely integrable if there is a family $A$ of lagrangian tori/abelian varieties which foliate $X$ such that, on a subset of full measure $\Sigma\subset X$, there is a diffeo/birational map $\delta:\Sigma\to A$ which identifies some iterate of $T$ with a relative translation $\tau:A\to A$.
A translation can be periodic, but generically it is quasi-periodic, meaning it winds up around the torus/abelian variety in a recurrent, equidistributed way.
3 Polygonal dynamic in $\mathbb{P}^{d}$
3.1 The collapse conjecture
A simple observation shows what are the potential collapse points. Generically, there will be only a finite number of candidates.
Lemma 3.1.
Proof.
@@ 3.1 The collapse conjecture
Let’s formulate the conjecture. It’s inspired by a question raised by Schwartz in his original paper [4, §4]. He stated in our private communication that no progress have been achieved since 1992.
@@
Let $k$ be an algebraicly closed, complete field for the metric coming from a non-trivial valuation. Let $T$ be a polygonal transformation taking place in $\mathbb{P}^{d}(k)$, such that the dynamic on the moduli space is discretely integrable.
Then for a generic choice of $n$-gon $P=((p_{1},\dots,p_{n}),M)$, there exists a matrix $\tilde{M}$ commuting with $M$, such that the vertices of $P$ collapse under the iteration of $T$ in the past/future towards $q_{-}$ and $q_{+}$, which are the repelling/attracting fixed points of $\tilde{M}$. Moreover, $q_{-}$ and $q_{+}$ depend continuously on $P$.
We only managed to prove the conjecture in the case of periodic motion on the moduli space. We know that such a periodic motion is possible for the pentagram map (see [27] and the link with Poncelet polygons) and for the staircase cross ratio dynamics (see subsection 6.2).
More generally, periodic motion over the moduli space occurs for polygons with $n$ is small enough, since their invariant tori/abelian varieties are points. As shown in [4], this happens when $n=5,6$ for the pentagram map over closed polygons. It also happens for closed $3$-gons in $\mathbb{P}^{1}$, since the moduli space is a point.
@@
Conjecture 3.2 holds for polygons whose dynamic is periodic on the moduli space.
@@
Let $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon and $N\in\mathbb{N}^{*}$ be such that $T^{N}([P])=[P]$. In the initial space, this means that there exists $\tilde{M}\in\mathbb{P}\mathrm{GL}_{d+1}$ such that $T^{N}(P)=\tilde{M}(P)$. By lemma 2.2, we can assume that $\tilde{M}$ generically has an attractive and a repelling fixed point. Since $T$ preserves the monodromy of $P$, we obtain:
$M=\tilde{M}M\tilde{M}^{-1}.$
Since $T$ commutes with the action of $\mathbb{P}\mathrm{GL}_{d+1}$, we have:
$T^{2N}(P)=T^{N}(\tilde{M}(P))=\tilde{M}(T^{N}(P))=\tilde{M}^{2}(P).$
By recurrence (that we can also apply backwards), we get:
$k\in\mathbb{Z},\penalty 10000\ T^{kN}(P)=\tilde{M}^{k}(P).$
We also have that for all $1\leq q<N,\penalty 10000\ k\in\mathbb{{Z}}$:
$T^{kN+q}(P)=T^{q}(T^{kN}(P))=T^{q}(\tilde{M}^{k}(P))=\tilde{M}^{k}(T^{q}(P)).$
We can assume that none of the $T^{q}(P)$ contains a vertex which is fixed by $\tilde{M}$. If we denote by $q_{\pm}$ the attractive/repelling fixed points of $\tilde{M}$, then:
$\lim_{k\to\pm\infty}T^{k}(P)=((q_{\pm},\dots,q_{\pm}),M)\qed$
Now the problem is: how to prove it for quasi-periodic motion, which is the truly generic case ? Since the motion is recurrent on the moduli space, one could try to mimic the case of periodic motion to say that some subsequence of the iterate “behaves asymptotically the same” as the iteration of a projective transformation.
The meaning of “behaving asymptotically the same” is yet unclear. One strategy could be to use generalized schwarzian derivatives (see [28]), but we haven’t succeeded.
3.2 Corollaries for closed polygons
The previous conjecture combined with lemma 3.1 allows us to characterize the collapse points as being the one of the monodromy of $P$. However, the original questions were formulated for closed polygons, and $\mathrm{Id}$ fixes every point, so we cannot predict the collapse points.
If the system we’re interested in admits a scaling symmetry [^9] Which should be the case, since this is how integrability is usually obtained., we can use the infinitesimal monodromy of a polygon to retrieve the “true fixed points” of the identity. The following corollary gives us a non-trivial analogue of lemma 3.1. Because of theorem 3.3, it is true for polygons with periodic motion on the moduli space, which we already gave some examples of.
@@ 3.2 Corollaries for closed polygons
Let $T$ be a polygonal transformation taking place in $\mathbb{P}^{d}(k)$ where $k$ is a complete non-trivially valued field, such that the dynamic is integrable on the moduli space and admits a scaling symmetry. Let $P$ be a generic closed polygon and $M^{\prime}$ be its infinitesimal monodromy.
Then the vertices of $P$ collapse under the iteration of $T$ in the past/future toward fixed point of $M^{\prime}$.
@@
Let $P$ be a generic closed polygon. According to conjecture 3.2, we know that the vertices of $P$ converge in the past/future towards $q_{-}$ and $q_{+}$, which are the repelling/attracting fixed points of $\tilde{M}$.
By applying the scaling symmetry, we obtain a family of polygons $P_{z}=((p_{1}(z),\dots,p_{n}(z)),M_{z})$. At least for each $z$ in a small neighborhood of $1$, we get by conjecture 3.2 that they collapse to some points $q_{-}(z)$, $q_{+}(z)$ which depend continuously on $z$. By genericity, we can assume that $M_{z}$ have distinct eigenvalues for $z\neq 1$ in this neighborhood. Since $\tilde{M}_{z}$ commutes with $M_{z}$, lemma 2.3 gives that $q_{-}(z)$, $q_{+}(z)$ are fixed points of $M_{z}$.
For $z\sim 1$ we have
$M_{z}=\mathrm{Id}+(z-1)M^{\prime}+o(z-1),$
so the fixed points of $M_{z}$ tend towards the ones of $M^{\prime}$. Since $q_{\pm}=q_{\pm}(1)$, we get the result. ∎
@@
Here are some details about its application to polygons with periodic motion on the moduli space, since the continuity part was not addressed in theorem 3.3.
The scaling symmetry commutes with the dynamic, so the dynamic of $P_{z}$ have the same periodicity on the moduli space for any $z$:
$T^{N}(P_{z})=\tilde{M}_{z}P_{z}.$
Since $\tilde{M}_{z}$ maps $(p_{1}(z),\dots,p_{d+1}(z))$ to $(p_{1}^{(N)}(z),\dots,p_{d+1}^{(N)}(z))$ and both $d+1$ tuples vary smoothly, we get by the fundamental theorem of projective geometry (see [14, Prop. 4.5.10]) that $\tilde{M}_{z}$ varies smoothly too. Hence its fixed points too. The nature of the repelling and attracting fixed points don’t change on a small enough neighborhood, so we have the result.
We also get a generalization of [6, Thm. 1.1].
Corollary 3.5.
Let $P$ be a closed $n$-gon in $\mathbb{P}^{d}$ and $(p_{i}=[p_{i}^{(0)}:\dots:p_{i}^{(d)}])_{i=1,\dots,n}$ be its vertices. Let $q=[q^{(0)}:\dots:q^{(d)}]$ be a collapse point predicted by corollary 3.4. Then $q^{(0)},\dots,q^{(d)}$ are contained in a field extension of $\mathbb{Q}\big((p_{i}^{(j)})_{i=1,\dots,n}^{j=0,\dots,d}\big)$ of degree at most $d+1$.

### v3

@@ 2.3 Linear scaling symmetry
Corollary 2.7 tells us that
$\mathrm{det}(M(t))=c_{\{1,\dots,n\}\times\{1,\dots,d\}}(t).$
Lemma 2.12 gives us the formula with the $R_{k}$’s. They are invariant by the dynamic since it preserves the monodromy, and by projective transformations because they are cross-ratios. ∎
For closed polygons, we can compute the higher derivatives of the monodromy.
Proposition 2.14.
Let $T$ be a polygonal transformation in $\mathbb{P}^{d}$ admitting a linear scaling symmetry. Let $P$ be a closed $n$-gon. Then, for any $k\in\mathbb{N}^{*}$, the $k$-th infinitesimal monodromy is
$M^{(k)}(1)=\sum_{1\leq i_{1}<\dots<i_{k}\leq n}S_{i_{1}}\dots S_{i_{k}},$
where the $S_{i}$’s are the matrices defined in equation 2. Again, it is invariant under the dynamic.
This provides a lot of preserved quantities, namely all the coefficients of the characteristic polynomial. We’ll give a formula for some of them in proposition 4.5.
Proof.
The computation is practically the same as the one done for theorem 4.2. By the linearity of the scaling symmetry, we know that $L_{i}^{\prime\prime}=0$ for any $i$. Hence, we get
$M^{(k)}(t)=\sum_{1\leq i_{1}<\dots<i_{k}\leq n}L_{1}(t)\dots L_{i_{1}-1}(t)L_{i_{1}}^{\prime}(t)L_{i_{1}+1}\dots L_{i_{k}-1}(t)L_{i_{k}}^{\prime}(t)L_{i_{k}+1}\dots L_{n}(t).$
we remark that for any $i<j$:
$L_{i+1}(t)\dots L_{j-1}(t)$ $=M_{i}(t)^{-1}M_{j-1}(t).$
When we evaluate at $t=1$, we have by the definition of $S_{i}$:
$M^{(k)}(1)$ $=\sum_{1\leq i_{1}<\dots<i_{k}\leq n}M_{i_{1}-1}(1)L_{i_{1}}^{\prime}(1)M_{i_{1}}^{-1}(1)\dots M_{i_{k}-1}(1)L_{i_{k}}^{\prime}(1)M_{i_{k}}^{-1}(1)$
$=\sum_{1\leq i_{1}<\dots<i_{k}\leq n}S_{i_{1}}\dots S_{i_{k}}.$
The invariance is again a consequence of the fact that the dynamic preserves the monodromy and commutes with the scaling action. ∎
3 Collapsing
We’ll only consider irreducible polygonal transformations, meaning that they cannot be decomposed in independent smaller systems. A counter-example would be any $2n$-gon $P$ in $\mathbb{P}^{2}$ composed of intertwining the vertices of two $n$-gons $P_{\text{even}}$ and $P_{odd}$ with the same monodromy, and applying the pentagram map respectively on them.
3.1 For general polygons
It has been known for a long time that the pentagram map acts trivially on the moduli spaces of closed pentagons and hexagons (see [schwartzPentagramMap1992, §2]). This phenomenon stops for polygons with at least seven sides. However there exist a family of closed polygons of any number of vertices (satisfying a geometric constrain), whose orbits on the moduli space are periodic ([schwartzPentagramIntegralsPoncelet2015], [izosimovPentagramMapPoncelet2022]).
This reasoning generalizes. The triviality of the action on the moduli spaces of $(d+3)$- and $(2d+2)$-gons holds for some pentagram maps in $\mathbb{P}^{d}$ (see [dirdakGALETRANSFORMPENTAGRAM]). Moreover, we prove in proposition 4.1 that any polygonal dynamic in $\mathbb{P}^{1}$ is trivial on the moduli space of triangles and quadrilaterals, and we construct in subsection 6.2 a family of closed $n$-gons (for any $n$) whose dynamic on the moduli space is periodic. More generally, for integrable systems, the invariant manifold of an $n$-gon is a point when $n$ is small enough, hence its dynamic is trivial on the moduli space.
Remark 9.
All of the previous examples involve closed polygons. But if the dynamic admits a scaling symmetry, it produces from a polygon $P$ a family of twisted polygons $(P_{t})$ whose dynamic is also periodic on the moduli space. This motivates the following theorem.
Theorem 3.1.
Let $F$ be an algebraicly closed field, complete for the metric coming from a non-trivial valuation. Let $T$ be a polygonal transformation taking place in $\mathbb{P}^{d}(F)$.
Then for almost every $n$-gon $P=((p_{1},\dots,p_{n}),M)$ whose dynamic is periodic on the moduli space, the vertices of $P$ collapse towards a fixed point of $M$.
The proof relies on the following lemmas, which we will prove later. The first one says that iterating a generic element of $\mathbb{P}\mathrm{GL}_{d+1}(F)$ yields a North-South dynamic, and the second one gives the potential collapsing points.
@@ 3.1 For general polygons
Let $F$ be an algebraicly closed field, complete for the metric coming from a non-trivial valuation
Then a generic element of $\mathbb{P}\mathrm{GL}_{d+1}(F)$ has one attractive and one repelling fixed point.
Lemma 3.3.
Now, let’s prove the theorem.
Proof of theorem 3.1.
Let $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon and $N\in\mathbb{N}^{*}$ be such that $T^{N}([P])=[P]$. In the initial space, this means that there exists $\tilde{M}\in\mathbb{P}\mathrm{GL}_{d+1}$ such that $T^{N}(P)=\tilde{M}\cdot P$. By lemma 3.2, we can assume that $\tilde{M}$ almost surely have an attractive and a repelling fixed point, respectively denoted by $q_{+}$ and $q_{-}$. Since $T$ commutes with the action of $\mathbb{P}\mathrm{GL}_{d+1}$, we have:
$T^{2N}(P)=T^{N}(\tilde{M}\cdot P)=\tilde{M}\cdot T^{N}(P)=\tilde{M}^{2}\cdot P.$
By recurrence (that we can also apply backwards), we get:
$l\in\mathbb{Z},\penalty 10000\ T^{lN}(P)=\tilde{M}^{l}\cdot P.$
We also have that for all $1\leq r<N$ and $l\in\mathbb{{Z}}$:
$T^{lN+r}(P)=T^{r}(T^{lN}(P))=T^{r}(\tilde{M}^{l}\cdot P)=\tilde{M}^{l}\cdot T^{r}(P).$
Then
$\lim_{l\to\pm\infty}T^{l}(P)=((q_{\pm},\dots,q_{\pm}),M),$
and by lemma 3.3 the collapse points have to be fixed points of $M$. ∎
We still have to prove the lemmas.
Proof of lemma 3.2.
Let $M\in\mathrm{GL}_{d+1}(F)$. The set of diagonalisable matrices of $\mathrm{GL}_{d+1}(F)$ is Zariski dense; indeed, $M$ is diagonalisable if its characteristic polynomial $\chi_{M}$ have distinct roots. This happens if its discriminant, which is a polynomial in the coefficients $\Delta(a_{1,1},\dots,a_{d+1,d+1})$, doesn’t vanish. Since this is not the zero polynomial, we obtain a Zariski-open set. It is not empty since it contains the identity.
@@
with $|\lambda_{1}|\leq\dots\leq|\lambda_{d+1}|$. Since the valuation is non-trivial, then out of genericity we can assume that the inequalities are in fact strict. In $\mathbb{P}\mathrm{GL}_{d+1}(F)$, we can take the representative with $\lambda_{d+1}=1$. So:
@@
Thus $M$ has an attractive fixed point in $\mathbb{P}^{d}(F)$, being $0:\dots:0:1]$ (the origin).
We can do similarly for $M^{-1}$ and take a representative with $\lambda_{1}=1$. This yields:
@@
Proof of lemma 3.3.
@@
It is known that some polygonal dynamics collapse for many polygons which aren’t periodic on the moduli space. The most famous example is the pentagram map on convex polygons [schwartzPentagramMap1992, Thm. 3.1]. It also happens for other pentagram maps in $\mathbb{P}^{2}$ [schwartzFlappingBirdsPentagram2026, Thm. 1.3] and in $\mathbb{P}^{d}$ (for any $d$) [yaoGlicksConjecturePoint2014a, Thm. 1.4]. Our computer simulations suggest that the collapsing almost always happens.
@@
Theorem 3.1 holds for almost every polygons, even if their dynamic is not periodic on the moduli space.
If the polygonal dynamic is integrable, then the motion on the moduli space is quasi-periodic. One could try to mimic the case of periodic motion and say that some subsequence of the iterate “behaves asymptotically the same” as the iteration of a projective transformation. [^10] This was already an idea from Schwartz in [schwartzPentagramMap1992, §4]. He stated in our private communication that no progress have been achieved since 1992. The meaning of “behaving asymptotically the same” is yet unclear. One strategy could be to use generalized schwarzian derivatives (see [ovsienkoProjectiveDifferentialGeometry2004]), but we haven’t succeeded.
Moreover, computer simulations seem to indicate that the conjecture is true for the projective heat map, which isn’t integrable. However, almost every projective classes of pentagons converge towards a fixed point ([schwartzProjectiveHeatMap2017, Cor. 1.7]) and it is conjecture to be true for any polygon, so they might also behave asymptotically like a projective transformation.
3.2 For closed polygons
By lemma 3.3, we know that the only potential collapsing points of a polygon are fixed by its monodromy. However, this is useless for closed polygons (whose monodromy is $\mathrm{Id}$). But if the dynamic admits a scaling symmetry, we can use the infinitesimal monodromy [^11] See definition 2.9. to retrieve the “true fixed points” of the identity.
@@ 3.2 For closed polygons
Let $P$ be a closed polygon satisfying the hypothesis of theorem 3.1. Moreover, suppose that the dynamic admits a scaling symmetry. Then, the collapse point $q$ is a fixed point of the infinitesimal monodromy $M^{\prime}(1)$.
Remark 10.
If $M^{\prime}(1)=0$, the same result holds by replacing it with the first non-vanishing derivative $M^{(k)}(1)$ defined in proposition 2.14.
This result should also be valid if the conjecture 3.4 is true, meaning that the collapse point is always a fixed point of the infinitesimal monodromy.
@@
Because of theorem 3.1, we know that the vertices of $P$ converge towards the point $q$. By applying the scaling symmetry, we obtain a family of polygons $P_{t}=((p_{1}(t),\dots,p_{n}(t)),M(t))$, where the dependence in $t$ is rational (in particular, smooth). Since the scaling symmetry commutes with the dynamic, every $P_{t}$ is periodic on the moduli space (with the same period $N$). Hence there is a family of projective transformations $(\tilde{M}(t))$ such that
$T^{N}(P_{t})=\tilde{M}(t)\cdot P_{t}.$
The function $\tilde{M}(t)$ is rational in $t$, since it is uniquely determined by the mapping
$(p_{1}(t),\dots,p_{d+2}(t))\mapsto(p_{1}^{(N)}(t),\dots,p_{d+2}^{(N)}(t)),$
and each of these vertices are rational in $t$. Since $\tilde{M}(1)$ generically have an attractive and a repulsive fixed point, it will still be the case for $t$ close enough to $1$. Hence the collapse point $q(t)$ of $P_{t}$ vary continuously. Moreover, we know by lemma 3.3 that the collapse point is a fixed point of $M(t)$. But since
$M(t)=\mathrm{Id}+(t-1)M^{\prime}+o(t-1),$
the fixed points of $M(t)$ tend towards the ones of $M^{\prime}$ when $t$ goes to $1$. Since $q=q(1)$, we get the result. ∎
We also get a generalization of [glickLimitPointPentagram2020, Thm. 1.1].
@@
Let $P$ be a closed $n$-gon in $\mathbb{P}^{d}$ with vertices
$(p_{i}=[p_{i,1}:\dots:p_{i,d+1}])_{i=1,\dots,n}.$
Let $q=[q_{1}:\dots:q_{d+1}]$ be the collapse point predicted by theorem 3.5. Then $q_{1},\dots,q_{d+1}$ are contained in a field extension of $\mathbb{Q}(p_{1,1},\dots,p_{n,d+1})$, of degree at most $d+1$.
@@
Since the scaling symmetry is an action of the algebraic group $\mathbb{G}_{m}$, the coefficients of the infinitesimal monodromy are rational functions in the $p_{i,j}$’s. Because of theorem 3.5, the collapse point $q=[q_{1}:\dots:q_{d+1}]$ is an eigenline of this $(d+1)\times(d+1)$ matrix. To find it, one needs to solve a linear system with $(d+1)$ equations involving the coefficients of the matrix, hence the result. ∎
Remark that our results don’t involve geometric considerations like convexity. They answer the question of Schwartz in [schwartzFlappingBirdsPentagram2026, §9.2] about the reappearance of Glick’s formula some generalized pentagram maps in $\mathbb{P}^{2}$.
@@
It would be interesting to consider similar statements in the setting of polygonal dynamics on $\mathbb{R}^{d}$ with the group $\mathrm{Isom}^{+}(\mathbb{R}^{d})$. For instance in $\mathbb{R}^{2}$, Benoist and Hulin prove in [benoistIterationPliagesQuadrilateres2004] that when $n=4$, the iteration of some element of the folding group is generically conjugated, on the moduli space, to a translation on an elliptic curve. This matches our definition 1.2 of integrability. The authors conclude that, on the initial space, the polygons drifts towards a point on the line at infinity, following some “drifting vector”. When seen on the projective plane, the fixed points of positive isometries are on the line at infinity, which is coherent with theorem 3.1.
4 Polygonal dynamics in $\mathbb{P}^{1}$
We have more effective statements for polygonal dynamics in $\mathbb{P}^{1}$. We apply them in the two last sections.
4.1 The infinitesimal monodromy and the limit point
Following definition 2.5, the cross-ratio coordinates of an $n$-gon $P$ in $\mathbb{P}^{1}$ is the $n$-periodic sequence
$c_{i}=[c_{i,1}:c_{i,2}],\quad i\in\mathbb{Z}.$
We normalize each $c_{i,2}$ to $1$, and we retrieve the cross-ratios already introduced in [arnoldCrossratioDynamicsIdeal2022]
$c_{i}=[p_{i},p_{i+1},p_{i-1},p_{i+2}]=\frac{p_{i}-p_{i-1}}{p_{i}-p_{i+2}}\frac{p_{i+1}-p_{i+2}}{p_{i+1}-p_{i-1}}\in\mathbb{P}^{1}\setminus\{0,\infty\}.$
Let’s start with a similar result than the one for the pentagram map on the moduli space of small closed polygons (see [schwartzPentagramMap1992]). This might generalize to higher dimensional projective spaces (see [dirdakGALETRANSFORMPENTAGRAM]).
Proposition 4.1.
Let $T$ be a polygonal dynamic in $\mathbb{P}^{1}$. Then we have $T=\mathrm{Id}$ on $\mathcal{P}_{3}$, and $T^{2}=\mathrm{Id}$ on $\mathcal{P}_{4}$.